# StableMaterials: Enhancing Diversity in Material Generation via Semi-Supervised Learning

### v2

@@ Abstract.
We detail the architecture and training process of StableMaterials, the integration of semi-supervised training within existing LDM frameworks, and show the advantages of our approach. Comparative evaluations with state-of-the-art methods show the effectiveness of StableMaterials, highlighting its potential applications in computer graphics and beyond.
StableMaterials is publicly available at https://gvecchio.com/stablematerials.
@@
journal: TOGccs: Computing methodologies Computer graphics
@@ 1. Introduction
Authoring of materials has been a long-standing challenge in computer graphics, requiring very specialized skills and a high level of expertise. To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [Deschaintre et al., 2018, 2019; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2023], or generation from a set of conditions [Guehl et al., 2020; Guo et al., 2020; Hu et al., 2022; Zhou et al., 2022; He et al., 2023; Vecchio et al., 2023, 2024]. While these approaches have reduced technical barriers to material creation, their effectiveness depends on the quality and diversity of training data, which can limit their use in real-world applications.
@@
To overcome these limitations, we introduce StableMaterials, an approach that takes advantage of semi-supervised adversarial training to: (1) include unannotated (non-PBR) samples in training, and (2) distill knowledge from a large-scale pretrained SDXL [Podell et al., 2023] model. In particular, we use a pretrained SDXL to generate unannotated texture samples from text prompts. However, since StableMaterials is trained to produce SVBRDF maps, we cannot perform direct supervision using the generated textures. To include these textures in the training of our methods, we learn a common latent representation between textures and materials; then, we complement the traditional supervised loss, with an unsupervised adversarial loss, forcing the model to also generate realistic maps for unannotated samples and close the gap between the two data distributions. In addition, drawing inspiration from the SDXL [Podell et al., 2023] architecture, we use a diffusion-based refinement model to enhance the visual quality of the samples and achieve high-resolution generation. We initially generate materials at the model base resolution of 512x512, and subsequently apply our refiner using SDEdit [Meng et al., 2021] and patched diffusion. This approach allows to achieve high resolution while constraining the patched generation, ensuring consistency and memory efficiency. Subsequently, we distill a latent consistency model [Song et al., 2023] that allows fast generation by reducing the number of inference steps to four steps per stage. However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [Vecchio et al., 2023] to achieve tileability. To solve this issue, we propose a novel features rolling technique, which shifts the tensor rolling from the diffusion step to the U-Net architecture by directly shifting the feature maps within each convolutional and attention layer.
We qualitatively evaluate our method and compare it with previous work, demonstrating the benefit of our semi-supervised training approach. In summary, we introduce StableMaterials a novel solution combining supervised and adversarial training to generate highly realistic material in scenarios where annotated data are scarce. The contributions of this work are as follows:
@@ 2. Related Work
Materials synthesis is an open challenge in computer graphics [Guarnera et al., 2016] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [Deschaintre et al., 2018; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Deschaintre et al., 2019; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2023]. Building on the success of generative models, several approaches to materials generation have emerged, including Guehl et al. [2020] which combines procedural structure generation with data-driven color synthesis; MaterialGAN [Guo et al., 2020], a generative network which produces realistic SVBRDF parameter maps using the latent features learned from a StyleGAN2 [Karras et al., 2020a]; Hu et al. [2022] which generates new materials transferring the micro and meso-structure of a texture to a set of input material maps; and TileGen [Zhou et al., 2022], a generative model capable of producing tileable materials, conditioned by an input pattern but limited to class-specific training. Recent appraoches have focused on leveraging the generative capabilities of diffusion models for materials generation. In particular, Vecchio et al. [2023] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image. MatFuse [Vecchio et al., 2024] extends generation control with multimodal conditioning, and enables editing ofexisting materials via ‘volumetric’ inpainting. MaterialPalette [Lopes et al., 2024] extends the capture of materials to pictures of real-world scenes by finetuning a LoRA [Hu et al., 2021] for each picture. Substance 3D Sampler [Adobe, 2024] recently introduced a pipeline to generate materials by first synthesizing a texture via text conditioning. However, these methods often lack diversity, struggling with complex material representation, or depend on image generation models, requiring additional steps to estimate the material parameters. StableMaterials overcomes these limitations by including a wider variety of unannotated material samples via a semi-supervised training and improves on inference time by distilling a latent consistency model.
@@ 3. Method
StableMaterials’ architecture, shown in Fig. 2, is based on MatFuse [Vecchio et al., 2024], which adapts the LDM paradigm [Rombach et al., 2022] to synthesize high-quality pixel-level reflectance properties for arbitrary materials. We modify the original MatFuse architecture, which employed a multi-encoder VAE to learn a disentangled latent representation of the maps, by fine-tuning a resource-efficient single-encoder compression model with the same latent properties. Additionally, we introduce a semi-supervised training strategy to distill knowledge from a large-scale SDXL model and improve generation diversity. We employ latent consistency distillation [Song et al., 2023] and a novel feature rolling technique for fast tileble generation, and achieve high resolution via a refiner model.
@@ 3.1. Material Representation
StableMaterials generates materials in the form of SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [Cook and Torrance, 1982; Karis, 2013], using a GGX [Walter et al., 2007] distribution function, as well as the material mesostructure. In particular, we generate base color, normal, height, roughness and metalness properties, with the roughness determining the BRDF specular lobe width and the metalness defining which area of the material represents a conductor.
@@ 3.2. Material Generation
The generative model consists of a compression VAE [Kingma and Welling, 2013] $\mathcal{E}$, encoding the material maps into a compact latent space, and a diffusion model [Rombach et al., 2022] $\epsilon_{\theta}$, modeling the distribution of these latent features.
@@
We first train a multi-encoder VAE, consisting of a set of encoders =$\mathcal{E}=\left\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{N}\right\}$ and a decoder $\mathcal{D}$, to encode $N$ material maps $M=\left\{\textbf{M}_{1},\textbf{M}_{2},\dots,\textbf{M}_{N}\right\}$ into their latent representation $z=concat\left(z_{1},z_{2},\dots,z_{N}\right)$, with $z_{i}=\mathcal{E}_{i}(\textbf{M}_{i})$ being the encoding of the $i^{th}$ map using the $i^{th}$ encoder $\mathcal{E}_{i}$.
The latent tensor $z$ is then decoded to reconstruct the input maps $\hat{M}=\mathcal{D}(z)$. Following [Vecchio et al., 2024], the VAE is trained using pixel-space $L_{2}$ loss, perceptual LPIPS loss [Zhang et al., 2021], a patch-based adversarial objective [Isola et al., 2017; Dosovitskiy and Brox, 2016], and a rendering loss [Deschaintre et al., 2018]. We regularize the VAE latent space using a Kullback-Leibler divergence penalty [Kingma and Welling, 2013; Rezende et al., 2014].
Following the training of the multiencoder VAE, we fine-tune a resource-efficient single encoder model. In particular, we freeze the decoder and train only the maps encoder $\mathcal{E}^{\prime}$ to compress the concatenated material maps to the same latent representation $z$. This encoder maintains the original’s disentangled latent representation and compression efficiency with fewer parameters. We compare the two models in Sec. 4.4.1.
Finally, we learn a common latent representation between texture images and materials, allowing us to incorporate these unannotated textures into the training process. To do so, we fine-tune an autoencoder model to compress a single texture (i.e.: the basecolor map) into $z$. Similarly to the single-encoder model, we keep the decoder frozen and train the encoder only. This network is used for semi-supervised training, allowing us to inject unannotated textures generated using SDXL in the training. We detail the use of this model in Sec. 3.3.
@@
We train a diffusion model on a compressed latent representation $z$ of the material. This model, based on the architecture of Rombach et al. [2022], uses a time-conditional U-Net [Ronneberger et al., 2015] to denoise the latent vectors $z$. During training, we generate noised latent vectors using a deterministic forward diffusion process $q\left(z_{t}|z_{t-1}\right)$, as defined in Ho et al. [2020], transforming them into an isotropic Gaussian distribution. The diffusion network $\epsilon_{\theta}$ learns the backward diffusion $q(z_{t-1}|z_{t})$ to denoise and reconstruct the original latent vector. The model training is described in Sec. 3.3
@@
StableMaterials enables control over the generation via high-level text or image prompts, descriptive of the global appearance of the material. Conditions are encoded into a one-dimensional feature vector using a pretrained CLIP [Radford et al., 2021] model. During training, we alternate between images and text prompts to enhance the robustness of the model. In particular, we use an ambient-lit rendering of the material as the image condition and the caption as text prompt. For materials without a caption, we use descriptive tags to build a text prompt. For each training batch, we select between image and text conditioning modes by randomly dropping one modality. A drop probability of 75% for text and 25% for image empirically proved to be the most effective.
@@ 3.3. Semi-Supervised Adversarial Distillation
To reduce the gap with image generation methods trained on large-scale datasets, we propose to distill knowledge from an SDXL model. However, direct distillation is impractical due to domain differences between textures, represented from a single image, and materials, represented by multiple maps; therefore, we propose a semi-supervised approach to include unannotated samples (i.e.: textures without explicit material properties representation) during training. In particular, we complement the supervised loss with an unsupervised adversarial loss through a discriminator on the latent space ($LD$). This adversarial loss forces the generator to synthesize materials that present features similar to ground-truth ones, while the discriminator learns to distinguish between the two data sources. The generator is trained using samples from the material dataset as well as unannotated textures, while the discriminator is trained using only annotated samples.
Our generator is optimized using a combination of supervised and adversarial losses. The supervised loss, $\mathcal{L}_{\text{sup}}$, is defined as:
(1) $\mathcal{L}_{\text{sup}}=\mathbb{E}_{t,z_{0},\epsilon}\left[\left\|\epsilon_{t},\epsilon_{\theta}\left(z_{t,\text{mat}},t\right)\right\|^{2}\right]+\alpha~{}\mathbb{E}_{t,z_{0},\epsilon}\left[\left\|\epsilon_{t},\epsilon_{\theta}\left(z_{t,\text{tex}},t\right)\right\|^{2}\right]$
comparing the model-estimated noise $\epsilon\theta(z_{t},t)$ with the actual noise $\epsilon_{t}$ introduced during the forward diffusion process [Ho et al., 2020]. Here, $z_{t,mat}$ and $z_{t,tex}$ are the noisy latent at timestep $t$ for the materials and textures samples respectively. The loss computed on unannotated texture samples is weighted using the hyperparameter $\alpha$, set to $0.15$ in our experiments.
The adversarial objective $\mathcal{L}_{\text{adv}}$ push the generator to close the distribution gap between materials and textures, producing outputs that are perceived as materials by the discriminator. It is computed on the denoised latents at timestep $t-1$, and is defined as:
(2) $\mathcal{L}_{\text{adv}}=-\mathbb{E}_{\mathbf{z}\sim p(\mathbf{z})}\left[LD(z_{t-1})\right]$
with $z_{t-1}=\text{concat}\left(z_{t-1,mat},z_{t-1,tex}\right)$ the concatenated denoised latents and $LD(\cdot)$ the output of the latent discriminator.
Following Sauer et al. [2021, 2023a], $LD$ is trained using a hinge loss function [Lim and Ye, 2017] computed between the noised latent embeddings of the materials $z_{t-1,\text{real}}$ encoded with the VAE, and $z_{t-1,\text{fake}}$ the denoised latents from the generator.
$LD$ shares the same architecture as the conditional diffusion U-Net encoder and is initialized with the weights of the diffusion U-Net. Functioning in a time-conditional manner, it incorporates the diffusion timestep and the CLIP embeddings from either text or image prompts, ensuring that evaluations are relevant to the desired material appearance.
Figure 3. Semi-Supervised training. Both the SDXL model and StableMaterials are prompted to generate the same material. The supervised $\mathcal{L}_{2}$ loss is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.

### v3

@@ Abstract.
We detail the architecture and training process of StableMaterials, the integration of semi-supervised training within existing LDM frameworks. Comparative evaluations with state-of-the-art methods show the effectiveness of StableMaterials, highlighting its potential applications in computer graphics and beyond. StableMaterials will be made publicly available.
@@
journal: TOGccs: Computing methodologies Appearance and texture representations
@@ 1. Introduction
Authoring of materials has been a long-standing challenge in computer graphics, requiring very specialized skills and a high level of expertise. To simplify the creation of materials for 3D applications, such as videogames, architecture design, simulation, media, and more, recent methods have tried to leverage learning-based approaches to capture materials from input images [Deschaintre et al., 2018, 2019; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2024a], or generation from a set of conditions [Guehl et al., 2020; Guo et al., 2020; Hu et al., 2022; Zhou et al., 2022; He et al., 2023; Vecchio et al., 2024a, b]. While these approaches have reduced technical barriers to material creation, their effectiveness depends on the quality and diversity of training data, which can limit their use in real-world applications.
@@
To overcome these limitations, we introduce StableMaterials, an approach that takes advantage of semi-supervised adversarial training to: (1) include unannotated (non-PBR) samples in training, and (2) distill knowledge from a large-scale pretrained SDXL [Podell et al., 2023] model. In particular, we use a pretrained SDXL to generate unannotated texture samples from text prompts. However, since StableMaterials is trained to produce SVBRDF maps, we cannot perform direct supervision using the generated textures. To include these textures in the training of our methods, we learn a common latent representation between textures and materials; then, we complement the traditional supervised loss, with an unsupervised adversarial loss, forcing the model to also generate realistic maps for unannotated samples and close the gap between the two data distributions. In addition, drawing inspiration from the SDXL [Podell et al., 2023] architecture, we use a diffusion-based refinement model to enhance the visual quality of the samples and achieve high-resolution generation. We initially generate materials at the model base resolution of 512x512, and subsequently apply our refiner using SDEdit [Meng et al., 2021] and patched diffusion. This approach allows to achieve high resolution while constraining the patched generation, ensuring consistency and memory efficiency. Subsequently, we distill a latent consistency model [Song et al., 2023] that allows fast generation by reducing the number of inference steps to four steps per stage. However, this comes at the cost of introducing visible seams when using approaches such as noise rolling [Vecchio et al., 2024a] to achieve tileability. To solve this issue, we propose a novel features rolling technique, which shifts the tensor rolling from the diffusion step to the U-Net architecture by directly shifting the feature maps within each convolutional and attention layer.
We evaluate our method qualitatively and quantitalively, and compare it with previous work, demonstrating the benefit of the semi-supervised training. In summary, we introduce StableMaterials a novel solution combining supervised and adversarial training to generate highly realistic material in scenarios where annotated data are scarce. The contributions of this work are as follows:
@@ 2. Related Work
Materials synthesis is an open challenge in computer graphics [Guarnera et al., 2016] with many recent data-driven approaches focusing on estimating SVBRDF maps from an image [Deschaintre et al., 2018; Li et al., 2017, 2018; Martin et al., 2022; Guo et al., 2021; Zhou and Kalantari, 2021; Vecchio et al., 2021; Deschaintre et al., 2019; Bi et al., 2020; Gao et al., 2019; Vecchio et al., 2024a]. Building on the success of generative models, several approaches to materials generation have emerged, including Guehl et al. [2020] which combines procedural structure generation with data-driven color synthesis; MaterialGAN [Guo et al., 2020], a generative network which produces realistic SVBRDF parameter maps using the latent features learned from a StyleGAN2 [Karras et al., 2020a]; Hu et al. [2022] which generates new materials transferring the micro and meso-structure of a texture to a set of input material maps; and TileGen [Zhou et al., 2022], a generative model capable of producing tileable materials, conditioned by an input pattern but limited to class-specific training. Recent appraoches have focused on leveraging the generative capabilities of diffusion models for materials generation. In particular, Vecchio et al. [2024a] introduced ControlMat, which relies on the MatGen diffusion backbone, to capture materials from an input image. MatFuse [Vecchio et al., 2024b] extends generation control with multimodal conditioning, and enables editing ofexisting materials via ‘volumetric’ inpainting. MaterialPalette [Lopes et al., 2024] extends the capture of materials to pictures of real-world scenes by finetuning a LoRA [Hu et al., 2021] for each picture. Substance 3D Sampler [Adobe, 2024] recently introduced a pipeline to generate materials by first synthesizing a texture via text conditioning. However, these methods often lack diversity, struggling with complex material representation, or depend on image generation models, requiring additional steps to estimate the material parameters. StableMaterials overcomes these limitations by including a wider variety of unannotated material samples via a semi-supervised training and improves on inference time by distilling a latent consistency model.
@@ 3. Method
StableMaterials builds on MatFuse [Vecchio et al., 2024b], which adapts the LDM paradigm [Rombach et al., 2022] to synthesize high-quality, pixel-level reflectance properties for arbitrary materials (Fig. 2). We replace MatFuse’s multi-encoder VAE architecture (used to learn a disentangled latent representation of material maps) with a more resource-efficient single-encoder model, fine-tuned to preserve the same latent properties. In addition, we introduce a semi-supervised training strategy that distills knowledge from a large-scale SDXL model to increase generation diversity. Our method leverages latent consistency distillation [Song et al., 2023] and a novel feature rolling technique for fast, tileable generation. A dedicated refiner model enables high-resolution outputs while preserving global consistency.
@@ 3.1. Material Representation
StableMaterials generates SVBRDF texture maps, representing a spatially varying Cook-Torrance microfacet model [Cook and Torrance, 1982; Karis, 2013], using a GGX [Walter et al., 2007] distribution function, as well as the mesostructure of the material. Specifically, the generated maps include base color, normal, height, roughness, and metalness, where roughness specifies the specular lobe width and metalness indicates conductor regions.
@@ 3.2. Material Generation
The generative model consists of a compression VAE [Kingma and Welling, 2013] $\mathcal{E}$, encoding the material maps into a latent space, and a diffusion model [Rombach et al., 2022] $\epsilon_{\theta}$, modeling the distribution of these latent features.
@@
We first train a multi-encoder VAE with encoders $\mathcal{E}={\mathcal{E}_{1},\dots,\mathcal{E}_{N}}$ and a decoder $\mathcal{D}$. Each encoder $\mathcal{E}_{i}$ encodes the $i^{th}$ map $\textbf{M}_{i}$ into a latent vector $z_{i}$, and the concatenated tensor $z=concat(z_{1},\dots,z_{N})$ is decoded to reconstruct the maps $\hat{M}=\mathcal{D}(z)$. Following [Vecchio et al., 2024b], training uses pixel-space $L_{2}$ loss, perceptual LPIPS [Zhang et al., 2021] loss, a patch-based adversarial loss [Isola et al., 2017; Dosovitskiy and Brox, 2016], and a rendering loss [Deschaintre et al., 2018], with a Kullback-Leibler penalty [Kingma and Welling, 2013; Rezende et al., 2014] to regularize the latent space.
Afterward, we fine-tune a single-encoder model. We freeze the original decoder and train only the encoder $\mathcal{E}^{\prime}$ to compress concatenated material maps into the same latent tensor $z$. This preserves the disentangled latent representation and maintains compression efficiency with fewer parameters, as shown in Sec. 4.4.1.
Lastly, we create a shared latent space for texture and materials. This allows us to have a common space that we can use to train our diffusion model. Specifically, we fine-tune an autoencoder that compresses a single texture (e.g., the base color) into $z$, again keeping the decoder frozen. This additional network is used for the semi-supervised training of the diffusion model, as described in Sec. 3.3.
@@
We train a diffusion model on the compressed latent representation $z$ of the material. This model, based on Rombach et al. [2022], uses a time-conditional U-Net [Ronneberger et al., 2015] to denoise the latent vectors $z$. During training, we generate noised latent vectors using a deterministic forward diffusion process $q\left(z_{t}|z_{t-1}\right)$, as defined in Ho et al. [2020], transforming them into an isotropic Gaussian distribution. The diffusion network $\epsilon_{\theta}$ learns the backward diffusion $q(z_{t-1}|z_{t})$ to denoise and reconstruct the original latent vector. The model training is described in Sec. 3.3
@@
StableMaterials allows flexible control over the generated material via text or image prompts describing high-level appearance. We encode the text or image condition using a pretrained CLIP model [Radford et al., 2021], which outputs a single feature vector. To make training robust, we alternate between image and text prompts. Specifically, we use (i) an ambient-lit rendering of the material as an image condition and (ii) a text caption for text prompts; when no caption is available, we generate short descriptive tags as an alternative. In each training batch, one modality is randomly dropped—text with 75% probability, image with 25%—to balance the two conditions.
@@ 3.3. Semi-Supervised Adversarial Distillation
Figure 3. Semi-Supervised training. Both the SDXL model and StableMaterials are prompted to generate the same material. The supervised $\mathcal{L}_{2}$ loss, between the estimated noise and the added noise, is complemented by an adversarial loss $\mathcal{L}_{\text{adv}}$ computed on the denoised latent from StableMaterials.
@@
To bridge the gap with image generation methods trained on large-scale datasets, we propose to distill knowledge from an SDXL model. However, direct distillation is impractical due to domain differences between textures, represented from a single image, and materials, represented by multiple maps; therefore, we propose a semi-supervised approach to include unannotated samples (i.e.: textures without explicit material properties) during training. Our method combines a supervised loss on annotated materials with an unsupervised adversarial loss that aligns the distribution of SDXL-generated textures with that of real materials in a shared latent space.
The supervised term ensures correct reconstruction of ground-truth material properties, while the adversarial term guides the model to learn from unannotated textures by treating them as partially labeled data. To accomplish this, we introduce a latent discriminator ($LD$) that distinguishes between real material features and generated ones, effectively forcing the generator to produce material-like features even from unannotated textures. Our training combines:
  1. (1)
    Supervised Loss ($\mathcal{L}{\text{sup}}$): Ensures reconstruction of material properties, maintaining physical plausibility.
  2. (2)
    Adversarial Loss ($\mathcal{L}{\text{adv}}$): Guides the generator to map both materials and textures to a shared feature distribution, enabling diversity while maintaining realism.
Our training strategy avoids mode collapse being primarely supervised, with the adversarial loss working as a distillation strategy.
Supervised Loss.
The supervised objective compares the denoiser’s prediction $\epsilon_{\theta}$ with the true noise $\epsilon_{t}$ introduced at each diffusion step $t$:
(1) $\mathcal{L}{\text{sup}}=\mathbb{E}_{t,z_{0},\epsilon}\Bigl{[}\bigl{|}\bigl{|}\epsilon_{t}-\epsilon_{\theta}\bigl{(}z_{t,\text{mat}},t\bigr{)}\bigr{|}\bigr{|}^{2}\Bigr{]}+\alpha~{}\mathbb{E}_{t,z_{0},\epsilon}\Bigl{[}\bigl{|}\bigl{|}\epsilon{t}-\epsilon_{\theta}\bigl{(}z_{t,\text{tex}},t\bigr{)}\bigr{|}\bigr{|}^{2}\Bigr{]}.$
Here, $z_{t,\text{mat}}$ and $z_{t,\text{tex}}$ are noisy latents at step $t$ for materials and textures, respectively. The hyperparameter $\alpha$ (set to $0.15$) controls the relative importance of unannotated texture samples.
Adversarial Loss.
In parallel, $\mathcal{L}{\text{adv}}$ encourages the model to treat textures and materials similarly, guiding the generator to produce material-like outputs even when starting from unannotated textures. We compute this loss on the denoised latents $z_{t-1}$:
(2) $\mathcal{L}{\text{adv}}=-\mathbb{E}{\mathbf{z}\sim p(\mathbf{z})}\Bigl{[}LD\bigl{(}z_{t-1}\bigr{)}\Bigr{]},$
where $z_{t-1}=\text{concat}\Bigl{(}z_{t-1,\text{mat}},z_{t-1,\text{tex}}\Bigr{)}$ is the concatenation of denoised latents for materials and textures, and $LD(\cdot)$ is the output of our latent discriminator. By using the discriminator to align the distributions of material and texture latents, this loss effectively distills knowledge from SDXL-generated textures while ensuring they conform to the features of real materials.
Latent Discriminator.
We follow Sauer et al. [2021, 2023a] and train $LD$ with a hinge loss [Lim and Ye, 2017], comparing real latent embeddings $z_{t-1,\text{real}}$ (encoded from the VAE) against fake latent embeddings $z_{t-1,\text{fake}}$ (denoised by the generator). The discriminator operates in a time-conditional fashion: it receives the same timestep $t$ and the CLIP embedding as the generator. Architecturally, $LD$ mirrors the U-Net encoder of the diffusion model and is initialized with the same weights, leveraging its understanding of the latent space to effectively guide the generator.
Unlike previous works that use adversarial distillation for fast generation [Sauer et al., 2023b, 2024], our approach bridges the domain gap between materials and textures. By training primarily with supervised terms while using adversarial guidance, the generator learns plausible material features even from unannotated textures. The adversarial component ensures SDXL-generated textures map to realistic material latents, effectively countering any shading artifacts and enriching the diversity of generated materials.